OpenStudy (anonymous):

Determine the quadratic function of f whose vertex is (1,-4) and passes through (2,-1) f(x)=?

OpenStudy (experimentx):

let f(x)=ax^2+bx+c the vertex is given by (-b/2a, f(-b/2a)) -b/2a = 1 ----------- 1 -4 = f(1) ----------- 2 -1 = f(2) ----------- 3 three equations three equations ... solve them to find the value of a,b,c

OpenStudy (experimentx):

or plug them into wolfram to get direct solution -b/2a = 1 ----------- 1 -4 = (1)^2a+(1)b+c ----------- 2 -1 = (2)^2a+(2)b+c ------------3